OpenStudy (anonymous):

Solve four and seven-eighths divided by two. sixteen thirty-ninths two and seven-sixteenths two and seven-eighths nine and three-fourths Question 2 (Multiple Choice Worth 1 points) [4.02] Add four and one-sixth plus three and one-third. seven and one-half seven and two-thirds eight and one-half eight and two-thirds Question 3 (Multiple Choice Worth 1 points) [4.02] Add three and one-eighth plus four and five-eighths. one and one-half seven and three-fourths seven and one-half one and three-fourths Question 4 (Multiple Choice Worth 1 points)

OpenStudy (anonymous):

I still need help

OpenStudy (anonymous):

where you stuck at? ^_^

OpenStudy (anonymous):

number 1

OpenStudy (anonymous):

(4 7/8) / 2 (turn to improper fraction) (39/8) / 2 (" flip " what you are dividing by, then multiply) 39/8 * 1/2 = 39/16 = 2 7/16

OpenStudy (anonymous):

4 1/6 + 3 1/3 set aside the whole numbers for now... 1/6 + 1/3 = 1/6 + 2/6 = 3/6 = 1/2 now deal with the whole numbers.. 4 + 3 = 7 put them together and you get : 7 1/2

OpenStudy (anonymous):

3 1/8 + 4 5/8 set aside the whole numbers.. 1/8 + 5/8 = 6/8 = 3/4 now add the whole numbers.. 3 + 4 = 7 put them together and you have : 7 3/4

OpenStudy (anonymous):

any questions ?